### v1

# Exploring the Impact of In-Browser Deep Learning Inference on Quality of User Experience and Performance

Abstract.
Deep Learning (DL) is increasingly being integrated into Web applications through a method known as ”in-browser inference”, where the DL processes occur directly within Web browsers. However, the actual performance of this method and its effect on user experience quality (QoE) is not well-understood. This gap in knowledge necessitates new forms of QoE measurement, going beyond traditional metrics such as page load time. To address this, we conducted the first extensive performance evaluation of in-browser inference. We introduced new metrics for this purpose: responsiveness, smoothness, and inference accuracy.
Our thorough study included 9 widely-used DL models and tested them across 50 popular PC Web browsers. The findings show a significant latency issue with in-browser inference: it’s on average 16.9 times slower on CPU and 4.9 times slower on GPU than native inference methods. Several factors contribute to this latency, including underused hardware instruction sets, inherent delays in the runtime environment, resource competition within the browser, and inefficiencies in software libraries and GPU abstractions.
Moreover, in-browser inference demands a lot of memory, sometimes up to 334.6 times more than the size of the DL models themselves. This excessive memory usage is partly due to suboptimal memory management. Additionally, we noticed that in-browser inference increases the time it takes for graphical user interface (GUI) components to load in web browsers by a significant 67.2%, which severely impacts the overall QoE for users of web applications that depend on this technology.
copyright: nonecopyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Human-centered computing Empirical studies in ubiquitous and mobile computingccs: Computing methodologies Artificial intelligence
1. Introduction
Deep Learning (DL) has been increasingly adopted in various Web applications by deploying DL models in Web browsers for inference tasks, i.e., in-browser inference (Ma et al., 2019; web, 2023b; fac, 2023; aig, 2023; off, 2023; tea, 2023). In this context, DL frameworks and models are encapsulated and delivered within Web pages without additional installations, and DL inference performs directly in Web browsers (Ma et al., 2019).
The proliferation of DL-powered Web applications has made in-browser inference performance a critical topic for the Web research community. However, existing research efforts primarily focus on DL inference performance in native and server environment (ten, 2023; Liu et al., 2022a; Zhang et al., 2022; Chen et al., 2022; Gao et al., 2020). In contrast, there is still a notable knowledge gap about DL in-browser inference performance in the literature. In addition, such performance is essential for the quality of experience (QoE) during Web browsing. Moreover, no previous work has considered the impact of in-browser inference on the QoE. Existing Web QoE measurement efforts primarily focus on the loading phase of Web pages such as page load time (spe, 2023a; Bocchi et al., 2016; Tian and Ma, 2019; PLT, 2023; Ling et al., 2018; Yeo et al., 2019), while, essentially, the QoE of in-browser inference is related to the whole process of user’s browsing experience.
To fill the knowledge gap, we present the first comprehensive measurement study of in-browser inference from two aspects: inference performance and QoE. Our extensive study involves 9 representative DL models across Web browsers of 50 popular PC devices. Furthermore, we use the most widely-adopted DL frameworks for deploying DL models in browsers, including TensorFlow.js (tfj, 2023b), ONNX Web Runtime (ort, 2023b).
For inference performance, we consider two critical metrics: latency and memory footprint. These metrics provide insights into the time overhead and memory requirements, respectively. We present a deep analysis of model-level/kernel-level latency and memory footprint, where a kernel is the basic unit of a DL model, such as a convolution operation. Furthermore, the inherent variability in device capacities motivates us to analyze the performance variance of in-browser inference across devices.
We observe a significant disparity in prediction latency between in-browser and native inference, averaging 16.9$\times$ on CPU and 30.6$\times$ on GPU. In the CPU case, this disparity arises primarily from the underutilization of Single-Instruction-Multiple-Data (SIMD) instructions, WebAssembly (Wasm) runtime-inherent inefficiency, as well as competition for resources with inherent browser tasks. On the GPU side, the inefficiency of software libraries on the Web, specifically WebGL (web, 2023a), compared to native CUDA (cud, 2023), contributes to the gap. Furthermore, our analysis reveals substantial performance variations across different devices, with disparities reaching up to 28.4$\times$ on CPU and 19.4$\times$ on GPU. These variations stem from the diverse capacities of devices and the hardware affinities of pre-designed DL kernels. Additionally, we identify ongoing concerns regarding memory footprint during in-browser inference. In certain scenarios, the memory footprint can escalate to 334.6$\times$ the model size during super-resolution model inference. These findings underscore the inefficiency of the in-browser inference on memory management when the intermediate data size increases.
For QoE, we are the first to measure it within the context of in-browser inference, proposing new metrics dedicated specifically to this endeavor beyond existing QoE metrics. Specifically, we quantitatively measure in-browser inference QoE from three aspects: responsiveness, smoothness, and inference accuracy. Responsiveness measures the speed at which user requests are successfully responded to. Smoothness evaluates the fluidity and seamless rendering of Web page content. Inference accuracy gauges the accuracy of DL models within Web applications. Our findings reveal an interesting observation: reducing inference latency does not necessarily equate to an improvement in QoE. For instance, while WebGL-based inference generally boasts faster execution time compared to Wasm, it also results in significant degradation of up to 62.7% in smoothness due to resource conflicts arising from the competition between inference and the web application’s reliance on GPU resources.
Based on the results, we provide a series of actionable implications for various stakeholders. (1) Web browser vendors can extend support for Wasm SIMD and streamline its integration with in-browser inference frameworks to avert potential performance degradation. (2) Web application developers can enhance the QoE by configuring inference to utilize the Wasm backend, albeit with a slight increase in inference latency. (3) DL framework vendors can employ efficient memory management techniques, such as model information-guided memory pool, to mitigate the memory overhead during inference.
2. Background
2.1. In-browser Inference
Web browsers have emerged as a pivotal platform for the widespread adoption of DL technologies (Chen et al., 2020; Ma et al., 2019). DL models find applications across various Web applications, enabling in-browser inference tasks that encompass a range of capabilities such as grammar checking for online document (off, 2023), background blurring for virtual meetings (tea, 2023), extended reality (Bi et al., 2023), and enhancing online chat interactions (web, 2023b).
In-browser inference is the process where Web applications execute DL inference directly within Web browsers. Figure 1 illustrates the workflow of this process, which comprises three primary stages: setup, warmup, and prediction. During the setup stage, Web applications download the necessary DL models and DL frameworks and then load them into Web browsers. In the warmup stage, browsers, assisted by DL frameworks, allocate the required memory for inference and initiate the first DL inference operation. Finally, in the prediction stage, DL models carry out subsequent inference tasks using the capabilities provided by DL frameworks.
Figure 1. Workflow of in-browser inference. “MB” denotes memory block.
2.2. DL Frameworks and Backends
Dedicated DL frameworks are developed to facilitate DL deployment in Web applications and in-browser inference. Among them, TensorFlow.js (TF.js) (tfj, 2023b) and ONNX Web Runtime (ORT.js) (ort, 2023b) are actively maintained and have garnered the most GitHub stars. TF.js is an offshoot of TensorFlow (TF) (Abadi et al., 2016) and maintains compatibility with it, while ORT.js is derived from ONNX Runtime (ORT) (ort, 2023b; onn, 2023b). In this study, we consider both of these frameworks due to their widespread adoption for in-browser inference in real-world Web applications (Ma et al., 2019; Guo et al., 2019; Chen et al., 2020).
Both TF.js and ORT.js support using backends for inference acceleration, including WebAssembly (Wasm) (Haas et al., 2017; was, 2023c) on CPU and WebGL (web, 2023c) on GPU. Wasm serves as a binary instruction format for a stack-based virtual machine within Web browsers. It serves the purpose of a portable compilation target for programming languages, facilitating their deployment on the Web. Wasm stands out for its efficiency, owing to its compact size and swift loading capabilities (was, 2023c). In Wasm, Single-Instruction-Multiple-Data (SIMD) and multithreading are available for computation efficiency. SIMD is a hardware architecture that processes multiple data elements in one instruction, greatly enhancing computational speed. Multithreading is a technique that allows a program to execute multiple tasks concurrently, making efficient use of system resources for improving performance. WebGL is a JavaScript library designed for rendering graphics directly within Web browsers, implementing an API that adheres to the OpenGL standards (ope, 2023). Developers can harness the power of WebGL by crafting WebGL shaders (web, 2023a) to expedite parallel computation processes.
Recently, a new backend called WebGPU has been introduced (web, 2023b). Nevertheless, the WebGPU standard has yet to attain official release status. WebGPU is still in its early and continually evolving phase, with no support for general models but only limited ones in TF.js. Therefore, we opt not to incorporate WebGPU into this paper.
3. Methodology
3.1. Experimental Setup
We start by outlining our choices for DL frameworks, backends, DL models, PC devices, and Web browser.
Selection of DL frameworks and backends. For in-browser inference, we opt for TF.js 4.2.0 and ORT.js 1.14.0 as our DL frameworks. Our selected backends include Wasm and WebGL, while we omit the pure JavaScript and WebGPU backends due to the lack of support for these options in ORT.js.
Selection of DL models. We select nine representative DL models for study, covering three common types of Web applications, i.e., image classification, object detection, and grammar checking. For each type, we select three widely adopted DL models. Specifically, for image classification, we select MobileNetV2 (denoted as M1; 14MB in size) (Sandler et al., 2018), ResNet50 (denoted as M2; 98MB in size) (He et al., 2016), and VGG16 (denoted as M3; 528MB in size) (Simonyan and Zisserman, 2014); for object detection, we select SSD-MobileNetV2 (denoted as M4; 26MB in size) (Liu et al., 2016), EfficientDet (denoted as M5; 23MB in size) (Tan et al., 2020), and YoloV5-Middle (denoted as M6; 83MB in size) (Redmon et al., 2016); for grammar checking, we select Bert-base (denoted as M7; 104MB in size) (Ahmad et al., 2019), MobileBert (denoted as M8; 95MB in size) (Sun et al., 2020), and small Bert (denoted as M9; 17MB in size) (Devlin et al., 2018). We use the pre-trained models (tfh, 2023; tf-, 2023) and convert them to the desired target format using official model converters (tfj, 2023a; onn, 2023a).
Selection of PC devices and Web browser. Our measurement involves 50 different PC devices, featuring a diverse array of CPU models, including both Intel CPUs and Apple Silicon. Within the Intel CPU category, we encompass core-i5 and core-i7 processors spanning from Skylake to Rocket Lake architectures (sky, 2023; roc, 2023). In the Apple Silicon category, we consider M1 and M2 CPU models. Moreover, these devices are equipped with a range of GPU models, covering both integrated and discrete graphics solutions. The discrete GPU options encompass Nvidia models such as GTX 980, 1060, and 2060, while integrated GPUs are embedded within the CPU and include variants such as Intel HD 530, 630, and 645, among others. For the Web browser, we employ Google Chrome, which stands as the most widely used browser, commanding a global market share of more than 63.5% as of September 2023 (bro, 2023).
3.2. Measurement of Inference Performance
First, we measure the performance of in-browser inference from two aspects: latency and memory footprint.
Latency. We measure the latency in both model-level and kernel-level. In terms of latency, the total inference latency consists of three types: setup latency, warmup latency, and prediction latency. (1) Model-level latency. We analyze the average latency gap and latency distribution among different DL frameworks, backends, and PC devices, and analyze the root cause. There are different acceleration techniques in the Wasm backend, i.e., SIMD and multithreading. The two DL frameworks that we adopt both support them for acceleration. SIMD processes multiple data within a single instruction and multithreading performs computation in parallel. To explore to what extent each acceleration technique reduces the latency, we measure the latency under different settings, i.e., whether each technique is enabled or not. (2) Kernel-level latency. A DL model can be represented as a series of interconnected kernels, such as matrix multiply and convolution. Understanding the performance of different kernels can help developers design kernel-specific optimization to improve the inference performance. For kernel-level latency, we measure the average kernel latency proportion on different backends. We do not consider the setup stage in this part because this stage does not involve kernel execution. In addition, we carry out a breakdown analysis to investigate how much SIMD and multithreading contribute to the acceleration in the Wasm backend.
Memory footprint. We next measure the memory footprint, i.e., the memory usage associated with in-browser inference. Memory is a critical metric both for the inference process (Wadhwani and Kojima, 2022; Molchanov et al., 2016) and Web applications as a whole (Qazi et al., 2020). To gain deep insights into the memory footprint of in-browser inference, we measure the memory footprint of each kernel and analyze the memory footprint growth. We omit the setup stage from our analysis since the inference execution does not take place during this phase. Similarly, we exclude the prediction stage from consideration because the DL framework retains allocated memory for operational efficiency, resulting in a constant memory footprint throughout this phase. As a result, we measure the memory footprint of only the warmup stage.
3.3. Definition of QoE Metrics Dedicated to In-browser Inference
Existing Web QoE metrics primarily focus on evaluating the speed at which a Web page becomes visually accessible during the page loading process (spe, 2023a; PLT, 2023; TTF, 2023). In contrast, the QoE associated with in-browser inference extends its influence across the entirety of a user’s browsing experience. Thus, we propose three new QoE metrics specialized for in-browser inference (including responsiveness, smoothness, and inference accuracy), and quantify it during model inference. For QoE measurement, we do not differentiate the three stages described in Figure 1, because QoE is an average metric that spans the entire Web application usage.
(1) $RunsPerMinute=\frac{\#Testcase}{Time_{Benchmark}}$
(2) $fps=\frac{\#Frames_{Rendered}}{Time_{Rendering}}$
(3) $InfAcc=\begin{cases}\frac{\#Frames_{correctlyClassified}}{\#Frames_{total}},&classification\\ \frac{1}{N}\sum_{i=1}^{N}\int_{0}^{1}PR_{i},&object\ detection\end{cases}$
$where\ PR_{i}\ is\ Precision\mbox{-}Recall\ curve\ of\ class\ i,\ and\ N\ is\ \#class.$
Responsiveness. Users prefer Web pages that respond to user requests quickly. The responsiveness reflects how quickly a Web page responds to user requests. It is quantified by the response time to user requests. We measure the responsiveness of Web pages based on Speedometer (spe, 2023b), a benchmark for Web application responsiveness. Speedometer simulates various Web activities, including 480 different tasks. As defined in Eq.1, we use the score “runs per minute” to quantify the responsiveness. The score represents the number of benchmark tests completed per unit time. We inject the inference code into the Speedometer Web page and perform in-browser inference during the benchmark test.
Smoothness. Users prefer Web pages that render smoothly without any lag. The smoothness reflects how timely the GUI content is rendered. It is quantified by the rendering/video frames per second (fps). As defined in Eq.2, we measure the fps of playing video and rendering Web pages. The fps stands for the average number of rendered frames per unit time during the entire testing phase. For vision tasks, we select YouTube as the target Web page for measurement; for the text task, we select Google Docs. The two Websites are among the most popular Websites for online videos and documents. YouTube provides continuous video streams and we can input text continuously into Google Docs. We measure the video fps in YouTube and rendering fps in Google Docs. We inject the inference code into each Web page. We run the inference for 10 minutes while using the Web page and use the averaged results as the final obtained fps.
Inference accuracy. Due to insufficient computation resource, inference accuracy might degrade inevitably. For example, video frames may be skipped and the application reuses previous results. Thus, as defined in Eq.3, we use inference accuracy, such as classification accuracy and detection mean average precision (mAP), to quantify the QoE. We use vision tasks, i.e., image classification and object detection, as the targets. We do not consider grammar checking in this part because, to our best knowledge, there lacks a benchmark dataset that contains both text and continuous timestamps. We use two benchmark datasets, including YouTube bounding boxes dataset (you, 2023a) for image classification and YouTube Face dataset (you, 2023b) for object detection. Both datasets contain continuous video frames and ground truth, as well as timestamps. Note that frames can use the previous inference results if they are not processed by DL models (Kang et al., 2017; Xu et al., 2020). We use the same models as before. We simulate in-browser inference with the fps obtained in the previous experiments.
4. Results of Inference Performance Measurement
Figure 2. ResNet50 average latency. “-I/-D” denotes integrated/discrete GPU. “-S/-T” denotes SIMD/multithreading.
4.1. Inference Latency Analysis
4.1.1. Model-level Latency
First, we compare the average inference latency in Web browsers and in native environments. We present the results of ResNet50 in Figure 2 and provide the complete results in the Appendix (Tables 4, 5, 6, and 7).
Average CPU latency. We observe that there is a latency gap between inference in the Wasm backend and in native. We analyze the gap of prediction latency, warmup latency, and setup latency. In this part, both SIMD and multithreading are enabled by default and we set 4 threads.
As for prediction latency, the in-browser prediction latency of TF.js is 3.7-18.4$\times$ higher than the latency of TF (Table 5 vs. 4); the gap is 2.1-36.4$\times$ between ORT.js and ORT (Table 6 vs. 4). The average gap of both frameworks is 16.9$\times$. The gap mainly comes from three aspects. First, advanced SIMD instruction sets are unavailable in browsers, like AVX2 and AVX512 on Intel CPU (AVX, 2023b, a). The SIMD instruction length in wasm is only 128-bit (was, 2023d), while the length is 256-bit in AVX2 and 512-bit in AVX512. This will introduce an up to 4$\times$ latency gap between in-browser and native inference. Besides instruction length, some advanced instructions, such as Fused-Multiply-Add (FMA) (fma, 2023), are not available in Wasm (was, 2023b). The absence of FMA instructions in Wasm could introduce a performance gap of up to 2x. Second, the Wasm VM introduces additional overhead, because it needs to interpret and execute the Wasm binary code, while native inference frameworks can directly run native binaries. In general, Wasm VM itself can lead up to 2.5$\times$ additional overhead (Jangda et al., 2019). Third, there are inherent threads in the Web page that introduce inevitable resource competition, such as the I/O thread and compositor thread (v8, 2023). Regarding the two in-browser inference frameworks, there is no consistent superiority of one framework over the other. We find that ORT.js exhibits lower prediction latency for ResNet50 when employing both multithreading and SIMD, while TF.js exhibits lower latency in inferring VGG16 and MobileNetV2.
As for warmup latency, the in-browser warmup latency of TF.js is 0.4-1.0$\times$ of the latency of TF (Table 5 vs. 4); while the gap is 1.0-20.5$\times$ for ORT.js and ORT (Table 6 vs. 4). TF suffers from a more severe cold-start issue across all models compared to TF.js. This is because of the large library size of TF (more than 980MB), while the library size of TF.js is only 3MB. Since memory allocation is not influenced by underlying hardware characteristics, the warmup latency gap between the browser and the native is not as significant as the gap in prediction latency. We find that warmup latency is 1.2-2.3$\times$ of the average prediction latency for TF.js (Table 5); while the gap is 1.0-1.8$\times$ for ORT.js (Table 6). Warmup additional overhead primarily arises from memory allocation in Wasm. After the warmup stage, the frameworks maintain the allocated memory for efficiency. The cold start issue is less severe in ORT.js compared to TF.js. This is primarily because ORT.js optimizes the model graph, reducing the frequency of memory allocations. Additionally, ORT.js leverages an Arena memory pool for further optimization (ort, 2023a, b).
As for setup latency, the in-browser inference setup latency of TF.js is 0.1-10.4$\times$ of the latency of TF (Table 4 vs. 5); while it is 0.9-2.4$\times$ for ORT.js and ORT (Table 4 vs. 6). It is worth noting that the setup stage of native inference only involves loading the model and the framework. In-browser inference may outperform native inference. One overhead of setup arises from loading the framework into memory. The native inference framework itself has a larger size. For example, simply importing TF in Python requires over 160MB memory, while the required files are only about 3MB for TF.js. On the other hand, native inference may also outperform in-browser inference because in-browser inference needs to download model files to the browser. The download latency depends on the network and model size. For example, in TF.js, downloading VGG16 requires downloading 528MB of data, which introduces high network transfer latency. Among the selected models, the model file size varies up to 31.1$\times$. The significant disparity in data transfer latency results in a wide range of setup latency. In most cases, native inference exhibits lower latency compared to in-browser inference.
Breakdown analysis for Wasm backend. We explore the influence of multithreading and SIMD on latency. We set 4 threads when enabling multithreading. The results of ResNet50 are illustrated in Figure 2 and the full results are presented in Appendix due to space limit (Table 5 and 6). The results show that SIMD reduces 49.1% and 50.7% average prediction latency for TF.js and ORT.js, respectively; while multithreading reduces 10.7% and 31.2% average prediction latency for TF.js and ORT.js, respectively. When enabling both techniques, the average prediction latency reduces by 64.7% and 63.4% for TF.js and ORT.js, respectively. For both frameworks, SIMD provides a greater speedup compared to multithreading. This is mainly because multithreading synchronization is implemented through message passing in Wasm, introducing additional overhead for thread management and synchronization (web, 2023a; was, 2023a). Both multithreading and SIMD can provide up to 4$\times$ acceleration (since 128-bit instruction length is 4$\times$ the length of int32/float32 data; single thread vs. 4 threads). This is because the model contains many memory-intensive kernels (37.7% on average across all models), and both frameworks provide limited optimization for these kernels.
Average GPU latency. We analyze the latency in WebGL and compare it with the latency in native and Wasm in the same aspects. Note that when comparing with native inference, we only consider discrete GPUs because the native DL framework does not support the integrated one.
As for prediction latency, TF.js is 2.5-268.7$\times$ slower than TF, and the gap between ORT.js and ORT is 11.2-17.9$\times$ (Table 4 vs. 7). The average gap for both frameworks is 30.6$\times$. The gap mainly arises from inefficient libraries and GPU abstraction. Native inference frameworks can utilize high-performance GPU libraries, such as CUDA (cud, 2023), to fully exploit the GPU’s parallelism capacity. CUDA’s low-level access to GPU allows developers to interact directly with the hardware, enabling fine-grained control and optimized program execution. However, in-browser inference frameworks rely on WebGL. WebGL is primarily designed for rendering tasks and requires using graphics-related mechanisms like explicit textures and frame buffers. Developers cannot access GPU directly through WebGL, making it inefficient for parallel computing tasks. We also find that WebGL exhibits lower latency compared with Wasm. For TF.js, inference in WebGL on integrated and discrete GPUs results in 1.4$\times$ and 2.7$\times$ average prediction latency reduction, respectively, compared to Wasm. For ORT.js, the reduction is 2.2$\times$ and 4.9$\times$, respectively. The speedup contributes to the GPU capacity, despite the limitations of WebGL in fully utilizing the hardware features.
As for warmup latency, TF.js is 5.8$\times$ slower than TF on average, and the gap between ORT.js and ORT is 186.7$\times$ (Table 4 vs. 7). The warmup latency of TF.js on discrete GPU is 64.6$\times$ longer than the prediction latency, while the gap is 26.8$\times$ on integrated GPU, and the gap for ORT.js is 48.1$\times$ and 25.1$\times$, respectively (Table 7). In-browser inference exhibits severe warmup latency in WebGL because the shader is compiled in this stage (web, 2023a).
As for setup latency, TF.js is 0.1-3.3$\times$ of TF, and the gap is 0.2-15.4$\times$ for ORT.js and ORT (Table 4 vs. 7). The reason is similar to that of Wasm. We also find that for in-browser inference, integrated GPU outperforms discrete GPU consistently. Specifically, the gap is 1.1-3.5$\times$ for TF.js and 1.5-2.8$\times$ for ORT.js (Table 7). This is mainly because of high data transfer and invoking overhead on discrete GPU (get, 2023; Web, 2023).
Figure 3. Prediction latency distribution. “M*” is model ID and is the same with §3.1
Prediction latency variance. Figure 3 shows the prediction latency variance. We exclude setup latency variance because network condition has a significant impact. We provide warmup latency variance results in the Appendix (Figure 6). The prediction latency varies differently among devices. Take ResNet50 in the Wasm backend of TF.js as an example, the latency quartiles are 257.6ms, 358.4ms, and 680.9ms. The prediction latency variance of all models reaches up to 28.4$\times$. Such variance mainly derives from hardware capacities. We observe that the benchmark score of the lowest-end CPU involved, Core i5 Skylake, is 4.1$\times$ lower than the highest-end CPU, Core i7 Rocket Lake (6,030 vs. 24,673) for one core (cpu, 2023). In the WebGL backend in TF.js, the prediction latency quartiles are 77.5ms, 124.3ms, and 225.3ms. The variance of all models reaches up to 19.4$\times$. Such variance also arises from the large device capacity gap. We observe that the score of the highest-end GPU in our dataset, Nvidia GTX 2060, is 207.4$\times$ better than the lowest-end Intel HD 530 according to the GPU benchmark (gpu, 2023) Since WebGL’s limited capacity to harness discrete GPUs, the variance is less pronounced compared to the differences in GPU capacity.
4.1.2. Kernel-level Latency
We explore the kernel-level inference performance of in-browser inference. Due to the immature support for profiling in ORT.js, we use TF.js for analysis. We provide warmup results in the Appendix (Table 8).
First, we explore which kernels dominate the prediction latency. Both multithreading and SIMD are enabled and we set four threads. Table 1 shows the latency proportion of top-5 kernels. We find that computation-intensive kernels predominantly contribute to the overall latency in both backends, although the number of memory-intensive kernels makes up 37.7% of the total. Specifically, Einsum and FusedConv2d contribute more than 70% of the total latency. We also find that the top-5 kernels are different in 2 backends due to different computation and memory-access patterns. In the Wasm backend, memory-intensive kernels introduce more latency. For example, GatherV2 accounts for the third most prediction latency in Wasm (12.5%), but it is not in the top-5 kernels in WebGL.
We explore the impact of SIMD and multithreading on the kernel, show the prediction latency results in Table 2, and provide warmup results in Appendix (Table 1). When enabling multithreading, we find that the prediction latency of memory-intensive kernels increases by up to 3.7%, and their latency proportion increases by up to 8.6%. This is because these memory-intensive kernels have limited support for multithreading. We dive into the kernel implementation in the source code and find that multithreading is only employed in some kernels, including Pad, among others. When enabling SIMD, we find that the latency proportion of computation-intensive kernels decreases by up to 13.0% and memory-intensive kernel latency changes negligibly. This is because Wasm has limited support for data-moving SIMD instruction, hindering the potential acceleration for memory-intensive kernels through SIMD.
These results suggest that memory-intensive kernels also require attention, as they can introduce high latency. The acceleration techniques have not yet been fully leveraged in memory-intensive kernels, emphasizing the need for framework vendors to focus on optimizing these kernels.
Finding 1: A large inference performance disparity exists between browser and native environments. The prediction latency gap in CPU is 16.9$\times$ on average. The shorter SIMD instructions and Wasm VM-inherent inefficiency are key reasons for the low performance of the Wasm backend. SIMD provides up to 50.7% latency reduction, and multithreading provides up to 31.2% latency reduction. WebGL significantly accelerates prediction by up to 4.9$\times$ compared to Wasm, but suffers from severe cold start issues, with a slowdown of up to 64.6$\times$, mainly due to in-browser shader compilation. The prediction latency spectrum across devices reaches up to 28.4$\times$ in Wasm and 19.4$\times$ in WebGL, which is mainly attributed to different hardware capacities. Memory-intensive kernels can contribute up to 12.5% of total latency.
Table 1. Top-5 kernels prediction latency proportion in TF.js. “Dw” denotes Depthwise. We enable SIMD and set 4 threads in Wasm. Each entry is the kernel prediction latency percentage out of the total latency of all models
Kernel Abbr.
Einsum K1
FusedConv2D K2
GatherV2 K3
Slice K4
FusedDwConv2D K5
Add K6
Multiply K7
FusedMatMul K8
Wasm WebGL-I WebGL-D
Abbr. Prop. Abbr. Prop. Abbr. Prop.
K1 41.4% K2 43.6% K1 48.4%
K2 34.6% K1 37.1% K2 23.7%
K3 12.5% K6 4.3% K6 7.9%
K4 2.8% K7 3.5% K7 3.4%
K5 1.9% K8 2.3% K8 2.9%
4.2. Memory Footprint Analysis
Table 2. Breakdown analysis of the impact of different acceleration approaches in Wasm backend, the framework is TF.js. “-S/T” indicates SIMD/multithreading. “Dw” indicates “Depthwise”. Each entry is the kernel prediction latency proportion. Cell color gray means the latency is a little longer than that of Wasm (1%¡ increment ¡5%); yellow means the latency is close to that of Wasm (difference ¡1%); blue means the latency is much less than that of Wasm (decrement ¿10%).
Wasm Wasm-S Wasm-T Wasm-S-T
FusedConv2D 47.6% 44.3% 47.4% 34.6%
EinSum 48.2% 46.5% 48.0% 41.4%
Gather 1.3% 3.9% 1.5% 12.5%
FusedMatMul 0.8% 1.8% 0.8% 1.6%
Slice 0.3% 1.1% 0.3% 2.8%
FusedDwConv2D 1.3% 1.2% 1.3% 1.9%
We find that the memory access pattern remains the same between the two backends and it is difficult to obtain WebGL memory usage in JavaScript. ORT.js implements and executes all logic in Wasm, hindering obtaining memory footprint in JavaScript. Owing to the above reasons, we only show the results in the Wasm backend of TF.js. SIMD and multithreading are enabled by default, but they do not affect the memory access. We measure the top-5 kernels that dominate the memory footprint during inference and illustrate the results in Figure 4(a). We find that memory-intensive kernels, such as Reshape, consume a significant amount of memory during inference, yet exhibit low latency.
To further understand the memory access patterns of different models, we illustrate the memory growth during warmup in Figure 4(b) because the framework only requires memory in this stage. We involve 3 super-resolution models in this experiment, i.e., ESRGAN (Wang et al., 2018), PGGAN (Karras et al., 2017) and CycleGAN (Zhu et al., 2017). These models are excluded from the latency analysis because ORT.js does not support them in both backends. We find that the curves of the models used in latency analysis show a similar pattern, i.e., initially increasing and then stabilizing. This is because the framework retains the allocated memory without releasing it back to the operating system. The memory footprint of the super-resolution model inference grows continuously. Take ESRGAN as an example, the peak memory footprint of ESRGAN inference bursts to 6,500MB, which is 334.6$\times$ of the model size, despite batch size being one and the model size being as small as 19MB. This is attributed to the expansion of intermediate data size, surpassing the size of previously allocated memory blocks, leading to additional requests for larger memory blocks from the operating system and a continuous increase in memory footprint. Such results emphasize the importance of efficient memory management design, as excessive memory footprint can hinder the ability for model inference.
Finding 2: Memory footprint can become a significant bottleneck during inference. For example, when performing super-resolution model inference like ESRGAN, the growing size of intermediate data can lead to a memory footprint surge of over 6,500MB, which is 334.6$\times$ of the model size. Inference of such models surpasses the capacity of most devices, necessitating the need for more effective memory management techniques in frameworks. The framework can utilize the model’s profile information and allocate a large enough memory block to meet peak memory footprint requirements.
Figure 4. Memory footprint analysis
5. Results of QoE Measurement
We select one model for each type of application and list their results in Table 3, and provide the full results in the Appendix (Table 9, 10, and 11) due to space limit.
5.1. Responsiveness Analysis
Without inference, the benchmark yields a mean score of 142.0, with a standard deviation of 2.3. When the DL is deployed, the mean score experiences a drop of up to 45.4% for both frameworks, while the standard deviation increases up to 5.7$\times$. The degradation primarily stems from resource occupancy. In terms of CPU and integrated GPU, the mean score difference is not significant. For both frameworks, the scores on the integrated GPU are less than 2% lower than the CPU. Since the benchmark contains tasks that run on both the CPU and GPU, the execution of the benchmark tasks is influenced by the DL inference in either backend. Moreover, the parallelism capacity of the integrated GPU is relatively weak, thereby impacting the final scores. In contrast, the discrete GPU achieved the highest scores in all settings due to its outstanding performance. We observed that the progress of benchmark tests slows down when inference and benchmark tasks run on the same hardware. These results indicate a trade-off between responsiveness and inference performance. Achieving better responsiveness and low inference latency simultaneously becomes challenging. Timely responsiveness requires more resources to be allocated for processing requests, leaving fewer resources available for inference tasks.
Finding 3: In-browser inference can significantly degrade the responsiveness by up to 45.4%. The reason is the computational resources required by inference compete with other critical components of the Web page.Allocating more CPU and GPU resources to minimize inference latency comes at the expense of responsiveness.
Table 3. Partial results of QoE. Each entry is in the format of “mean (standard deviation)”. “/” indicates the model is not supported. “-D/-I” denotes the discrete/integrated GPU. “M*” denotes model ID and is consistent with §3.1.
Rresponsiveness Smoothness Inference accuracy
M1 M4 M7 M1 M4 M7 M1 M4
Baseline 142.0   (2.3) 142.0   (2.3) 142.0   (2.3) 30.0 (0.00) 30.0 (0.00) 60.0 (0.00) 17.2% (0.00) 1.00 (0.00)
TF.js-wasm 92.7   (7.5) 112.0   (7.0) / 23.8 (1.83) 23.9 (1.65) / 13.7% (0.53×$\times 10^{-2}$) 0.78 (2.91×$\times 10^{-2}$)
ORT.js-wasm 92.0   (5.0) 108.0 (11.0) 92.1 (6.6) 24.1 (1.79) 23.8 (1.79) 54.6 (4.07) 13.7% (0.50×$\times 10^{-2}$) 0.78 (3.10×$\times 10^{-2}$)
TF.js-WebGL-I 87.6 (12.0) 113.0 (13.0) 89.3 (9.5) 11.2 (0.77) 12.7 (0.82) 26.8 (2.10) 8.7% (0.33×$\times 10^{-2}$) 0.76 (2.86×$\times 10^{-2}$)
TF.js-WebGL-D 118.1   (2.9) 127.4   (7.4) 113.5 (3.1) 27.2 (1.29) 28.3 (1.71) 52.4 (2.44) 14.9% (0.54×$\times 10^{-2}$) 0.95 (3.26×$\times 10^{-2}$)
ORT.js-WebGL-I 89.7   (7.7) / / 17.0 (1.11) / / 10.5% (0.39×$\times 10^{-2}$) /
ORT.js-WebGL-D 123.1   (5.1) / / 26.4 (1.39) / / 15.1% (0.57×$\times 10^{-2}$) /
5.2. Smoothness Analysis
Figure 5 illustrates the fps trace during SSD-MobileNetV2 inference. Without inference, the average Web page rendering fps on the device is 60, while the average video playing fps is 30. When DL is deployed, the fps value inevitably drops (up to 43.4%). The integrated GPU exhibits the poorest performance in terms of smoothness, while the performance of the CPU is remarkably good, approaching that of the discrete GPU. Specifically, when performing DL model inference on the CPU, the fps experiences a relatively smaller drop, averaging 36.2% for both frameworks, compared to performance inference on the integrated GPU. Discrete GPU experiences 8.1% less fps degradation than the CPU and 44.3% less fps degradation than the integrated GPU. Furthermore, we observe that fps fluctuations are more severe when inferring on the integrated GPU, as depicted by the red dashed line in Figure 5. In Figure 5, the fps amplitude reaches 4, which is 31.5% of the average value (12.7). The fps amplitude can reach up to 37.3% of the average fps for all models. The reason is that rendering and video playing rely on GPU resources, leading to inevitable resource competition when performing inference in WebGL. The best performance on discrete GPU is attributed to its high capacity. Moreover, in WebGL, the discrete GPU is not fully harnessed, allowing rendering and video-playing tasks to utilize the remaining capabilities.
While performing inference on integrated GPU offers lower prediction latency, it may lead to fps degradation, because of its relatively lower capacity. On the contrary, performance inference on the CPU is slower but leads to less fps compromise since it alleviates resource competition. These results indicate the existence of a trade-off between smoothness and inference performance. In terms of smoothness, the priority of using integrated GPU for inference is lower than other hardware.
Finding 4: In-browser inference impacts smoothness, with fps dropping by up to 62.7%. This is due to resource competition. The resource competition can lead up to 37.3% fps fluctuations. Achieving both smoothness and inference performance is a trade-off. Using the Wasm backend for inference can lead to better smoothness compared to the WebGL backend on devices without a powerful discrete GPU.
Figure 5. Fps trace during SSD-MobileNetV2 inference.
5.3. Inference Accuracy Analysis
In the ideal scenario where all images can be processed by the DL model, the classification accuracy is 17.2% and the detection mAP is 1.0. This lower classification accuracy is attributed to the dataset containing low-resolution images, some of which are even smaller than the common 32$\times$32 pixel size used in many classification tasks. The high detection mAP is mainly because the object region typically covers a significant portion of the image, and the dataset exclusively consists of images with a single class (face). This characteristic is well-suited for in-browser detection tasks, like background blurring, where humans are the predominant subject. Additionally, the dataset we use comprises videos sourced from real-world video streams.
When performing inference in browser, we observe that the classification accuracy drops by up to 9.1%, and the detection mAP drops by up to 24%. Similar to §5.2, the performance of the integrated GPU remains inferior, while the CPU shows better performance. The reduced video fps and increased inference latency result in the DL model being unable to process all frames. As a result, some frames have to reuse the inference results from the previous frame, leading to a decrease in inference accuracy (Canel et al., 2019; Kang et al., 2017; Chen et al., 2015). In summary, in-browser inference leads to inference accuracy degradation. The constrained computational resources and reduced fps pose challenges for the DL model to process all inputs effectively. Performing inference in a slower backend may bring better inference accuracy.
Finding 5: In-browser inference can degrade the inference accuracy. We observe a decrease of up to 9.1% in classification accuracy and 24% in detection mAP. When performing inference in Wasm, applications could achieve higher inference accuracy at the cost of a minor latency increase, compared with inference in WebGL on integrated GPU. The results show that reducing inference latency does not necessarily equate to an improvement in QoE.
6. Implications and Suggestions
This section discusses a series of actionable implications based on our findings.
For Web browser vendors: (1) Extend Wasm SIMD instructions (Finding 1). Limited SIMD instructions in Wasm hinder the potential optimization. Browser vendors can consider extending the SIMD instructions with more features, such as FMA and data-moving instructions. This enhancement can empower the in-browser inference to leverage advanced hardware capabilities better. (2) Support loading pre-compiled WebGL binaries (Finding 1). Browser vendors can consider extending the WebGL API to accommodate the loading of pre-compiled binaries, to avoid time-consuming in-browser WebGL shader compilation. This improvement can not only benefit the in-browser inference but also enhance the performance of other Web applications.
For DL framework vendors: (1) Compile WebGL binaries ahead-of-time (Finding 1). In-browser compilation introduces significant warmup overhead to in-browser inference. Given the limited and hardware-agnostic WebGL compilation targets (e.g., Direct3D on Windows (dir, 2023) and Metal on OS X (met, 2023)), pre-compiling WebGL binaries to these targets and loading them on demand can mitigate the issue. This suggestion complements and synergizes with the second suggestion for Web browser vendors. (2) Focus more on memory-intensive kernels (Finding 1). Memory-intensive kernels can introduce high latency, underscoring the importance of paying attention to these kernels. For example, framework vendors can make efforts to fully utilize multithreading in the implementation of these kernels. (3) Manage memory through an efficient memory pool (Finding 2). When the size of intermediate data surpasses the allocated memory block size, frameworks need to require more memory. To mitigate the problem, memory pool techniques, optimized by profiling inference process or DL model, can be employed for efficient memory management (Wang et al., 2022; Zhou et al., 2023). For example, the memory access pattern can be saved during model conversion and the inference framework can allocate all memory directly during the setup stage. The size of the memory block should be able to meet the peak memory usage requirements. In this way, data can be accommodated within the block, redundant memory allocation.
For Web application developers: (1) Incorporate DL models with varying capacities (Finding 1). Web application developers can encapsulate multiple models within the application, each with different sizes and capacities, such as MobileNet (Sandler et al., 2018; Howard et al., 2019) and ResNet (He et al., 2016) catering to image classification, and download the suitable model according to hardware information. Web application developers can dynamically select the optimal model without introducing additional packaging and setup overhead. (2) Prioritize responsiveness (Finding 3). In Chrome’s architecture, the main thread of each Web page manages the GUI and interacts with other threads. Starting a separate thread for inference or intermittent inference can prevent interference with the responsiveness of the Web page. The two techniques are mutually complementary to one another. (3) Enhance smooth rendering (Finding 4). Since rendering is typically GPU-dependent, using the Wasm backend for inference could guarantee smooth rendering. While Wasm has slightly higher inference latency than WebGL, it circumvents rendering resource competition, especially for devices with only integrated GPUs. This can introduce a significant improvement in Web page smoothness at the cost of a little inference latency. (4) Guarantee inference accuracy (Finding 5). For the video applications we used, frame filtering techniques (Li et al., 2020) can effectively enhance inference accuracy. Additionally, utilizing Wasm as the inference backend is a practical choice, aligning with the previous suggestions and the findings.
7. Related Work
DL deployment. DL deployment focuses on fitting models into the different platforms after DL model is well tested (Mittal et al., 2016; Xu et al., 2018; Guo et al., 2019; Chen et al., 2018). Many frameworks facilitate the deployment in the cloud, like TensorFlow (Abadi et al., 2016) for servers, TFLite (tfl, 2023) for Android, and Core ML (cor, 2023) for iOS. Many studies are proposed to promote DL deployment (Xu et al., 2019; Chen et al., 2021, 2020; Cummaudo et al., 2020; Guo et al., 2019). For example, Xu et al. (Xu et al., 2019) presented the first measurement study on large-scale Android applications to characterize how DL is adopted in mobile devices. Guo et al. (Guo et al., 2019) investigated the performance gap when the pre-trained DL models are migrated from PC to mobile devices. Cummaudo et al. (Cummaudo et al., 2020) analyzed pain points that developers face when deploying DL to servers.
DL performance. DL performance is critical to software development and deployment, as well as to end users’ QoE. For example, high accuracy and low inference latency are desired (Xu et al., 2018; Bonawitz et al., 2019). To characterize the DL performance, many engineering and research efforts have been presented (Ma et al., 2019; ten, 2023; Ha and Zhang, 2019; Cao et al., 2022; Chen et al., 2022; Pham et al., 2020; Du et al., 2019; Gao et al., 2020). For example, TensorBoard (ten, 2023) is a visualization toolkit for visualizing DNN structures, tracking training processes, and profiling programs; DeepPerf (Ha and Zhang, 2019) is an end-to-end DL-based solution that can predict the performance of a new DL task through a pre-trained model; DNNMem (Gao et al., 2020) estimates the memory consumption of both the DL model and the DL framework systematically and accurately. Additionally, Ma et al. (Ma et al., 2019) measured the in-browser inference latency at model level for manually constructed models. In comparison, our study aims to anatomize in-browser inference of real-world DL models at both model level and kernel level, as well as the memory footprint and user QoE of in-browser inference.
Web browsing performance. Many metrics have been proposed to quantify the Web browsing performance, such as Page Load Time (PLT) (PLT, 2023), Time-to-First-Byte (TTFB) (TTF, 2023), Speed Index (spe, 2023a), etc. Several efforts have been presented to explore and optimize the performance. Lighthouse (lig, 2023) is a tool to analyze the quality of Web pages and is integrated into Chrome. WProf (Wang et al., 2013) is proposed to profile the PLT, which is a key performance metric that many techniques aim to reduce (Kelton et al., 2017; Mardani et al., 2021). SipLoader (Liu et al., 2022b) is a Speed Index oriented page scheduler. Enrico et. al proposed a comprehensive study to measure the QoE of web users (Bocchi et al., 2016). However, they mainly focus on the loading phase, and the performance after the loading stage is not well-studied.
8. Conclusion
We anatomized the in-browser inference performance and its impact on QoE. Results show that: (1) the latency gap between in-browser inference and native inference reaches up to 16.9$\times$ on CPU and 30.6$\times$ on GPU. The discrepancy in CPU is mainly attributed to different SIMD instruction sets and Wasm runtime-inherent inefficiency, as well as resource competition within the browser; the discrepancy in GPU is mainly attributed to inefficient software libraries and GPU abstractions. (2) Memory can be the bottleneck. When inferring super-resolution models, the memory requirement can reach up to 334.6$\times$ the size of the models themselves. (3) QoE will degrade due to inevitable resource competition. Inferring models on CPU may bring better QoE due to the tread-off between inference latency and QoE.
Appendix A Appendix: Additional Results
Table 4. Latency (ms) of different models running in native. The GPU used is the discrete one.
TensorFlow ONNX Runtime
CPU GPU CPU GPU
setup warmup prediction setup warmup prediction setup warmup prediction setup warmup prediction
(M1) MobilenetV2 1351.5 463.5 17.6 2069.9 2074.8 7.9 140.0 6.8 2.9 1394.2 1271.8 2.8
(M2) Resnet50 6407.3 997.4 82.2 6223.4 1000.4 11.2 488.0 25.4 23.4 130.7 187.7 5.6
(M3) VGG16 873.2 2308.4 153.6 3050.8 681.2 3.4 2057.3 90.4 101.8 839.8 126.7 8.2
(M4) SSD-MobilenetV2 9359.6 2571.7 43.3 10031.3 5128.5 65.0 1245.9 41.7 29.4 1562.8 730.2 65.0
(M5) EfficientDet 27722.3 4823.6 482.6 27107.2 6522.6 228.1 2056.7 478.7 337.1 2522.4 9124.3 274.4
(M6) Yolo5-Middle 11719.9 1209.3 84.3 10599.7 1562.8 39.5 517.7 49.8 41.7 249.3 288.4 8.0
(M7) Bert-base 5153.0 856.4 11.7 4938.3 1196.8 19.1 614.2 39.6 31.0 453.8 18.0 16.0
(M8) MobileBert 29766.9 3526.9 18.5 30586.4 3269.2 42.5 970.4 77.4 53.2 1215.7 61.8 56.4
(M9) Bert-Small 2044.5 168.3 3.6 2037.6 109.7 5.0 212.6 4.6 2.2 126.7 6.0 3.8
Table 5. Latency (ms) of supported models in TF.js in Wasm backend. “-SIMD/thread” indicates enabled technique.
Wasm Wasm-thread Wasm-SIMD Wasm-thread-SIMD
setup warmup prediction setup warmup prediction setup warmup prediction setup warmup prediction
(M1) MobilenetV2 381.5 320.9 235.0 335.2 284.0 222.2 400.3 194.5 116.6 254.3 161.5 89.2
(M2) Resnet50 2362.6 1984.3 1710.3 2109.7 1724.7 1416.1 2485.1 899.7 539.5 1665.2 647.3 300.8
(M3) VGG16 10283.0 8303.8 6069.9 10265.0 7597.0 5334.9 10747.3 5482.7 3060.7 9111.0 2402.5 1021.5
(M4) SSD-MobilenetV2 574.6 1386.6 1160.6 575.2 1174.6 1064.6 608.2 1024.1 833.4 474.8 962.4 795.0
Table 6. Latency (ms) of models in ORT.js Wasm backend. “-SIMD/thread” indicates enabled technique.
Wasm Wasm-thread Wasm-SIMD Wasm-thread-SIMD
setup warmup prediction setup warmup prediction setup warmup prediction setup warmup prediction
(M1) MobilenetV2 581.4 148.1 120.6 1306.6 216.9 122.4 1252.1 162.8 106.5 331.0 139.6 105.5
(M2) Resnet50 539.3 1299.5 1275.3 699.1 395.1 394.7 2412.3 411.9 358.1 1219.3 306.9 172.3
(M3) VGG16 7545.3 9024.7 8275.6 4437.5 8147.2 8140.1 2695.2 3164.6 2947.2 2423.2 1353.6 1334.1
(M4) SSD-MobilenetV2 1120.5 287.6 279.4 1338.1 197.3 123.9 1171.8 174.5 126.1 1467.8 84.5 81.4
(M5) EfficientDet 1581.2 2665.4 2648.6 1786.4 1251.6 1282.5 1606.1 1245.8 1228.8 1889.6 926.0 852.4
(M6) Yolo5-Middle 427.6 1105.2 1092.7 450.0 365.9 362.9 408.1 406.2 386.4 457.6 145.8 145.2
(M7) Bert-base 744.6 915.2 910.2 702.4 285.8 279.6 681.8 246.2 247.2 574.8 92.8 81.6
(M8) MobileBert 873.5 855.1 849.8 935.8 310.1 291.1 865.0 239.3 236.2 945.0 120.2 113.6
(M9) Bert-Small 155.2 21.2 21.8 161.3 9.2 7.8 158.1 7.1 6.7 166.8 4.6 4.7
Table 7. Latency (ms) of models in WebGL backend. “-I/-D” denotes integrated/discrete GPU.
ORT.js-WebGL-I ORT.js-WebGL-D TF.js-WebGL-I TF.js-WebGL-D
setup warmup prediction setup warmup prediction setup warmup prediction setup warmup prediction
(M1) MobilenetV2 210.7 2094.4 33.1 306.6 2866.8 31.4 209.5 2900.7 28.7 530.9 5751.0 20.1
(M2) Resnet50 720.6 1599.6 183.0 2011.8 2810.6 78.2 1635.7 2814.2 248.3 2482.7 5443.6 79.8
(M3) VGG16 733.5 1793.5 545.6 2086.3 2538.4 146.4 9109.0 3414.5 1748.7 10090.0 4719.0 913.9
(M4) SSD-MobilenetV2 / / / / / / 503.1 7047.6 1039.0 1125.2 9097.9 616.6
(M5) EfficientDet-D1 / / / / / / 734.8 9235.6 2736.1 2542.3 11127.3 1342.7
(M6) Yolo5-Middle / / / / / / 1739.7 5796.1 99.4 2054.3 10873.6 82.9
(M7) Bert-base / / / / / / 1902.7 3126.1 682.0 3412.0 3128.2 133.4
(M8) MobileBert / / / / / / 2837.3 3169.5 943.1 3538.0 2196.4 400.5
(M9) Bert-Small / / / / / / 664.0 2089.0 41.1 839.3 2273.4 58.3
Table 8. Kernel warmup latency proportion in TF.js. “Dw”, “-S/T” , and cell color indicates the same as Table 2 in the draft. Each entry is the kernel warmup latency proportion out of the total warmup latency.
Kernel Wasm Wasm-S Wasm-T Wasm-S-T Kernel WebGL-I WebGL-D
Einsum 33.6% 34.5% 36.0% 34.4% Einsum 38.1% 36.7%
FusedConv2D 55.1% 52.3% 53.4% 52.1% FusedConv2D 49.5% 52.5%
GatherV2 2.6% 3.1% 2.4% 4.5% Add 1.8% 2.6%
Slice 0.6% 1.2% 2.9% 3.8% Multiply 2.4% 3.4%
FusedMatMul 2.4% 2.1% 1.4% 1.3% FusedMatMul 1.9% 2.0%
FusedDwConv2D 2.1% 1.3% 1.1% 1.1% FusedDwConv2D 3.1% 2.8%
Figure 6. Warmup latency distribution. “M*” is model ID and is the same with §3.1.
Table 9. Responsiveness of Web page. The format of each entry, “/”, “-D/-I”, and “M*” keeps the same as Table 3.
M1 M2 M3 M4 M5 M6 M7 M8 M9
TF.js-Wasm 92.7   (7.5) 89.5 (11.0) 91.2 (7.0) 112.0   (7.0) / / / / /
ORT.js-Wasm 92.0   (5.0) 94.5   (9.2) 88.4 (9.8) 108.0 (11.0) 92.6 (12.0) 91.0   (7.0) 92.1 (6.6) 115.0 (12.2) 83.7   (7.0)
TF.js-WebGL-I 87.6 (12.0) 84.2   (7.5) 94.2 (8.3) 113.0 (13.0) 100.4  (9.3) 97.4 (11.2) 89.3 (9.5) 106.0   (7.7) 83.1 (12.5)
TF.js-WebGL-D 118.1   (2.9) 119.2   (3.8) 114.5 (3.6) 127.4   (7.4) 121.3  (6.3) 112.0   (2.4) 113.5 (3.1) 120.8   (3.6) 95.5   (8.6)
ORT.js-WebGL-I 89.7   (7.7) 88.2   (8.4) 96.2 (4.4) / / / / / /
ORT.js-WebGL-D 123.1   (5.1) 119.7   (6.2) 125 (4.0) / / / / / /
Table 10. Smoothness of Web page. The format of each entry, “/”, “-D/-I”, and “M*” keeps the same as Table 3.
M1 M2 M3 M4 M5 M6 M7 M8 M9
TF.js-wasm 23.8 (1.83) 22.9 (1.79) 22.4 (1.61) 23.9 (1.65) / / / / /
ORT.js-wasm 24.1 (1.79) 23.1 (1.49) 24.5 (1.71) 23.8 (1.79) 21.7 (1.66) 22.4 (1.66) 54.6 (4.07) 53.7 (3.71) 52.7 (3.51)
TF.js-WebGL-I 11.2 (0.77) 11.7 (0.85) 12.9 (1.01) 12.7 (0.82) 11.1 (0.84) 11.5 (0.79) 26.8 (2.10) 27.2 (1.82) 23.5 (1.73)
TF.js-WebGL-D 27.2 (1.29) 26.4 (1.18) 28.1 (1.51) 28.3 (1.71) 27.9 (1.26) 28.1 (1.49) 52.4 (2.44) 51.5 (3.19) 52.0 (2.85)
ORT.js-WebGL-I 17.0 (1.11) 16.8 (1.11) 17.1 (1.31) / / / / / /
ORT.js-WebDL-D 26.4 (1.39) 23.0 (1.08) 24.5 (1.14) / / / / / /
Table 11. Inference accuracy on certain applications. The format of each entry, “/”, “-D/-I”, and “M*” keeps the same as Table 3.
M1 M2 M3 M4 M5 M6
TF.js-Wasm 13.7% (0.53×$\times 10^{-2}$) 13.5% (0.52×$\times 10^{-2}$) 8.8% (0.33×$\times 10^{-2}$) 0.78 (2.91×$\times 10^{-2}$) / /
ORT.js-Wasm 13.7% (0.50×$\times 10^{-2}$) 13.6% (0.52×$\times 10^{-2}$) 8.0% (0.30×$\times 10^{-2}$) 0.78 (3.10×$\times 10^{-2}$) 0.77 (2.84×$\times 10^{-2}$) 0.79 (3.00×$\times 10^{-2}$)
TF.js-WebGL-I 8.7% (0.33×$\times 10^{-2}$) 8.8% (0.32×$\times 10^{-2}$) 3.9% (0.14×$\times 10^{-2}$) 0.76 (2.86×$\times 10^{-2}$) 0.76 (2.93×$\times 10^{-2}$) 0.78 (2.83×$\times 10^{-2}$)
TF.js-WebGL-D 14.9% (0.54×$\times 10^{-2}$) 14.2% (0.52×$\times 10^{-2}$) 11.1% (0.44×$\times 10^{-2}$) 0.95 (3.26×$\times 10^{-2}$) 0.94 (3.41×$\times 10^{-2}$) 0.96 (3.59×$\times 10^{-2}$)
ORT.js-WebGL-I 10.5% (0.39×$\times 10^{-2}$) 11.1% (0.45×$\times 10^{-2}$) 7.9% (0.30×$\times 10^{-2}$) / / /
ORT.js-WebGL-D 15.1% (0.57×$\times 10^{-2}$) 15.1% (0.56×$\times 10^{-2}$) 13.5% (0.54×$\times 10^{-2}$) / / /